# The spherical maximal operators on hyperbolic spaces

### v2

@@ Abstract.
In this article we investigate $L^{p}$ boundedness of the spherical maximal operator $\mathfrak{m}^{\alpha}$ of (complex) order $\alpha$ on the $n$-dimensional hyperbolic space $\mathbb{H}^{n}$, which was introduced and studied by Kohen [13]. We prove that when $n\geq 2$, for $\alpha\in\mathbb{R}$ and $1<p<\infty$, if
$\|\mathfrak{m}^{\alpha}(f)\|_{L^{p}(\mathbb{H}^{n})}\leq C\|f\|_{L^{p}(\mathbb{H}^{n})},$
then we must have $\alpha>1-n+n/p$ for $1<p\leq 2$; or $\alpha\geq\max\{1/p-(n-1)/2,(1-n)/p\}$ for $2<p<\infty$. Furthermore, we improve the result of Kohen [13, Theorem 3] by showing that $\mathfrak{m}^{\alpha}$ is bounded on $L^{p}(\mathbb{H}^{n})$ provided that $\mathop{\mathrm{Re}}\alpha>\max\{{(2-n)/p}-{1/(pp_{n})},\ {(2-n)/p}-(p-2)/[pp_{n}(p_{n}-2)]\}$ for $2\leq p\leq\infty$, with $p_{n}=2(n+1)/(n-1)$ for $n\geq 3$ and $p_{n}=4$ for $n=2$.
@@ 1.1. Background
provided that $\alpha>1-n+n/p$ when $1<p\leq 2;$ or $\alpha>(2-n)/p$ when $2\leq p\leq\infty.$ The above maximal theorem tells us that when $\alpha=0$ and $n\geq 3$, the operator ${\mathfrak{M}}$ is bounded on $L^{p}({\mathbb{R}^{n}})$ for $p>n/(n-1)$. This range of $p$ is sharp. As is pointed out in [24, 25], no such result can hold for $p\leq n/(n-1)$ if $n\geq 2$.
@@
[…]
@@
Note that in [17], Nowak, Roncal and Szarek [17] found sharp conditions for the operator ${\mathfrak{M}}^{\alpha}$ on radial functions to be bounded on $L^{p}_{\rm rad}({\mathbb{R}}^{n})$. They proved that for $n\geq 2$ and $\alpha>(1-n)/2$, ${\mathfrak{M}}^{\alpha}$ is bounded on $L^{p}_{\rm rad}({\mathbb{R}}^{n})$ if and only if $\alpha>1-n+n/p$ if $1<p\leq 2$; $\alpha>1/p+(1-n)/2$ if $2\leq p\leq 2n/(n-1)$; and $\alpha\geq(1-n)/p$ if $p>2n/(n-1)$.
@@ 1.2. Main results
The purpose of this article is to investigate the spherical maximal operators of (complex) order $\alpha$ on the $n$-dimensional hyperbolic space $\mathbb{H}^{n}$, which was introduced and studied by Kohen ([13]). To do this, we denote the Minkowski metric on ${\mathbb{R}^{n+1}}$ by
$[z,w]=z_{0}w_{0}-z_{1}w_{1}-\cdots-z_{n}w_{n},\ \ \ \ z=(z_{0},\dots,z_{n}),w=(w_{0},\dots,w_{n})\in{\mathbb{R}^{n+1}}.$
@@
and we equip this space with the Riemannian metric induced by the above Minkowski metric [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ]. This Riemannian metric induces in turn a measure, by which we denote $\mathrm{d}z$. Define the spherical maximal operators by
@@
A natural question is for which $\alpha$ the spherical maximal opeators ${\mathfrak{m}}^{\alpha}$ are bounded on $L^{p}$. In [13, Theorem 3], Kohen showed that for $n\geq 2$,
@@
or
[…]
where the two dimensional case was left open since the necessary condition $p>2$ for $\alpha=0$ eliminates the use of the $L^{2}$ argument based on Plancherel’s theorem. It was not until 2000 that Ionescu [10] showed that ${\mathfrak{m}}^{0}$ is bounded on $L^{p}(\mathbb{H}^{2})$ for $2<p\leq\infty$ by making use of the local smoothing estimates for Fourier integral operators ([20]). The above two admissible relations for $\alpha$ and $p$ are summarized when $n\geq 3$ in Figure 1 below, where the relations (1.10) and (1.11) correspond to the dashed segments $AB$ and $OB$, respectively.
[…]
Figure 1. The $(1/p,\alpha)$-plot for $n\geq 3$. Kohen’s result [13, Theorem 3] indicates that $\mathfrak{m}^{\alpha}$ is bounded on $L^{p}(\mathbb{H}^{n})$ if $(1/p,\alpha)$ is strictly above the dashed folded segments $OBA$. We extend the range of $(1/p,\alpha)$ in Kohen’s result for $2<p<\infty$ to the range strictly above the dotted folded segments $OCB$. In addition, we showed that $\mathfrak{m}^{\alpha}$ is bounded on $L^{p}(\mathbb{H}^{n})$ only if $(1/p,\alpha)$ is on or above the solid folded segments $ODB$ for $2<p<\infty$, or strictly above the dashed segment $BA$ for $1<p\leq 2$.
@@
[…]
@@
Theorems  1.1 and  1.2, can be illustrated using Figure 1 above. The relations (1.12) and (1.13) in Theorem 1.1 correspond to the dashed segment $AB$ and the solid folded segments $ODB$, respectively. Theorem 1.2 indicates that compared to (1.11), there is an $1/(pp_{n})$-downwards extension of the range of $\mathop{\mathrm{Re}}\alpha$ so that (1.9) is valid for $p\geq p_{n}$, and there is a $(p-2)/[pp_{n}(p_{n}-2)]$-improvement for $2<p<p_{n}$. This is indicated in Figure 1 above by the dotted folded segments $OCB$.
@@ 2. Fourier analysis on the hyperbolic space
$[z,w]=z_{0}w_{0}-z_{1}w_{1}-\cdots-z_{n}w_{n},\ \ \ \ \ z=(z_{0},\dots,z_{n}),\ w=(w_{0},\dots,w_{n})\in{\mathbb{R}^{n+1}}.$
@@
and equip this space with the Riemannian metric induced by the Minkowski metric [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ]. This Riemannian metric induces in turn a measure, by which we denote $\mathrm{d}z$. We will further distinguish the point ${\bf 0}=(1,0,\dots,0)$.
@@
(notice that $[z,b(\omega)]>0$ for $z\in\mathbb{H}^{n}$). They satisfy
@@
$\left|{\mathrm{d}^{k}\over\mathrm{d}\lambda^{k}}(\lambda^{-1}({\mathcal{H}}(\lambda))^{-1})\right|\leq C_{\alpha}\,(1+|\lambda|)^{(n-1)/2-1-k}.$
@@
$\left|{\mathrm{d}^{k}\over\mathrm{d}\lambda^{k}}(\lambda\,{\mathcal{H}}^{\alpha}(\lambda))\right|\leq C_{\alpha}\,(1+|\lambda|)^{1-(n-1)/2-\mathop{\mathrm{Re}}\alpha-k}.$
@@
This is a consequence of Stirling’s formula ([27, p. 151, Example. (ii)]) ∎
@@
In [13, Theorem 1], Kohen showed that for $\alpha\in\mathbb{C}$ with $\mathop{\mathrm{Re}}\alpha>(1-n)/2$ one can write
@@
×$\times\left({\mathfrak{R}}^{\alpha}(\lambda,t)+t^{\alpha+(n-1)/2}\sum_{j=0}^{N}t^{2j}b^{\alpha}_{j}(t)\,\mathfrak{O}_{\alpha+(n-2)/2+j}(\lambda t)\right),$
@@
×$\times F\!\left(-\alpha-{n-3\over 2},\alpha+{n-1\over 2};1-\mathrm{i}\lambda;{-1\over\mathrm{e}^{2t}-1}\right)$
@@
$\|\mathfrak{m}^{\alpha}(g_{j})\|_{p}$$\geq$ ($\left(\int_{\{z=(z_{0},z_{1},z^{\prime\prime})\in\mathbb{H}^{n}:\rho(z)\sim 1,|z^{\prime\prime}|\leq c_{2}2^{-j/2}\}}|M^{\alpha}_{\rho(z)}(g_{j})(z)|^{p}\,\mathrm{d}z\right)^{1/p}$
@@
Now we consider $z\in B_{\varepsilon/100}(\mathbf{0})$. Then it’s obvious that $\mathop{\mathrm{supp}}h_{\varepsilon}$ is contained in the interior of $B_{1}(z)$, which indicates that if $w\in\mathop{\mathrm{supp}}h_{\varepsilon}$, then we have $01-d(z,w)>$ 0. Thus in view of (1.7) for such $z$ we can write
@@
where $a^{\alpha}_{2}(\lambda,t)\in S^{0}_{\lambda}$ and ${\mathcal{H}}^{\alpha}$ is as in (2.4). Set $\phi_{s}\in C^{\infty}_{c}(\mathbb{R})$, $0\leq\phi_{s}\leq 1$, $\phi_{s}\equiv 1$ on $[0,s+1]$ and $\phi_{s}\equiv 0$ outside $(-1,s+2)$. The convolution kernel $K^{\alpha}_{t}(r)$ of $M^{\alpha}_{t}$ can be written as the following:
@@
×$\times\int_{\mathbb{R}}\beta_{j}(\lambda)\,{\mathcal{H}}^{\alpha}(\lambda)\left({\mathrm{e}^{\mathrm{i}\lambda(t+r)}\over{\mathcal{H}}(-\lambda)}a_{2}^{0}(\lambda,r)+{\mathrm{e}^{\mathrm{i}\lambda(t-r)}\over{\mathcal{H}}(\lambda)}a_{2}^{0}(-\lambda,r)\right)a_{2}^{\alpha}(\lambda,t)\,\mathrm{d}\lambda$
@@
×$\times\int_{0}^{\infty}\beta_{j}(\lambda)\left(\mathrm{e}^{\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(\lambda)\,a^{\alpha}_{2}(\lambda,t)+\mathrm{e}^{-\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(-\lambda)\,a^{\alpha}_{2}(-\lambda,t)\right)\varphi_{\lambda}(r)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda.$
@@
Now we assume $j\geq 1$. We take an even cutoff function $\gamma\in C^{\infty}_{c}(\mathbb{R})$ with $\gamma\equiv 1$ on $[-1,1]$ and $\gamma\equiv 0$ outside $(-2,2)$. By Lemma 2.4(ii), we write
@@
×$\times\int_{0}^{\infty}\beta_{j}(\lambda)\,\gamma(\lambda r)\left(\mathrm{e}^{\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(\lambda)\,a^{\alpha}_{2}(\lambda,t)+\mathrm{e}^{-\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(-\lambda)\,a^{\alpha}_{2}(-\lambda,t)\right)\varphi_{\lambda}(r)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda,$
@@
×$\times\int_{\mathbb{R}}\beta_{j}(\lambda)\,(1-\gamma(\lambda r))\left(\mathrm{e}^{\mathrm{i}\lambda(t+r)}a^{0,N}_{1}(\lambda,r)+\mathrm{e}^{\mathrm{i}\lambda(t-r)}a^{0,N}_{1}(-\lambda,r)\right){\mathcal{H}}^{\alpha}(\lambda)\,a^{\alpha}_{2}(\lambda,t)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda,$
@@
×$\times\int_{0}^{\infty}\beta_{j}(\lambda)\,(1-\gamma(\lambda r))\left(\mathrm{e}^{\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(\lambda)\,a^{\alpha}_{2}(\lambda,t)+\mathrm{e}^{-\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(-\lambda)\,a^{\alpha}_{2}(-\lambda,t)\right)E^{\alpha,N}(\lambda,r)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda.$
@@
$\sup_{z\in\mathbb{H}^{n}}\left\|\sup_{t\in[T,T+1]}|\mathscr{E}^{\alpha}_{t,j}(d(z,\cdot))|\right\|_{1}+\sup_{w\in\mathbb{H}^{n}}\left\|\sup_{t\in[T,T+1]}|\mathscr{E}^{\alpha}_{t,j}(d(\cdot,w))|\right\|_{1}\leq C_{\alpha,\sigma}\mathrm{e}^{-(n-1)t/2}$
@@
Let us now prove (4.22). To do this, we define another cutoff function $\widetilde{\phi}_{s}\in C_{c}^{\infty}(\mathbb{R})$ by $\widetilde{\phi}_{s}\equiv\phi_{s}$ on $[s,\infty)$ and $\widetilde{\phi}_{s}\equiv 0$ on $(-\infty,s-1]$. Note that in view of (4.12), we can write
@@
In view of (4.25), we set the function $\Phi((v,u),(v^{\prime},u^{\prime}))$ in (4.4) to be $d((v,u),(v^{\prime},u^{\prime}-T_{0}))$ and the symbol $a(\lambda,t,(v,u),(v^{\prime},u^{\prime}))$ in (4.4) to be $b^{\alpha}(\lambda,t,d((v,u),(v^{\prime},u^{\prime}-T_{0})))$. Notice that the problem in $[(v,u),(v^{\prime},u^{\prime})]$-coordinates is no longer degenerate. Indeed, on the one hand, by (4.29) the Monge-Ampere matrix of $\Phi$ turns into
@@
which is (4.2). Here $\nabla$ is the gradient and $|\cdot|_{g}$ is the norm both induced by the Riemannian metric on $\mathbb{H}^{n}$. Also, in view of (4.26) and (4.29), $b^{\alpha}(\lambda,t,d((v,u),(v^{\prime},u^{\prime}-T_{0})))$ satisfies (4.3) with $m$ replaced by $-\mathop{\mathrm{Re}}\alpha$. Hence by (4.7) in Proposition 4.3, we can write
@@
Now we consider (4.33) for the case $0<t\leq 1$. Let $\gamma$ be the cutoff function in the proof of Lemma 4.6. We define another cutoff function $\widetilde{\gamma}\in C_{c}^{\infty}(\mathbb{R}_{+})$ with $\widetilde{\gamma}\equiv 1$ on $[1/2,\infty)$ and $\widetilde{\gamma}\equiv 0$ on $[0,1/4]$. Let $K^{\alpha}_{t}(r)$ be the convolution kernel of $M^{\alpha}_{t}$. Note that in view of (1.7), $\mathop{\mathrm{supp}}K^{\alpha}_{t}(r)\subset\{r:0\leq r\leq t\}$. So we can write
@@
Acknowledgements: The authors were supported by National Key R&\&&D Program of China 2022YFA1005700. P. Chen was supported by NNSF of China 12171489. The authors thank X. Chen, N.J. Liu, L. Roncal, L. Song and H.W. Zhang for helpful discussions and suggestions.

### v3

@@ Abstract.
In this article we investigate $L^{p}$ boundedness of the spherical maximal operator $\mathfrak{m}^{\alpha}$ of (complex) order $\alpha$ on the $n$-dimensional hyperbolic space $\mathbb{H}^{n}$, which was introduced and studied by El Kohen. We prove that when $n\geq 2$, for $\alpha\in\mathbb{R}$ and $1<p<\infty$, if $\mathfrak{m}^{\alpha}$ is bounded on $L^{p}(\mathbb{H}^{n})$, then we must have $\alpha>1-n+n/p$ for $1<p\leq 2$; or $\alpha\geq\max\{1/p-(n-1)/2,(1-n)/p\}$ for $2<p<\infty$. Furthermore, we improve El Kohen’s result [J. Operator Theory 3 (1980)] on $L^{p}$ boundedness of $\mathfrak{m}^{\alpha}$ by showing that $\mathfrak{m}^{\alpha}$ is bounded on $L^{p}(\mathbb{H}^{n})$ provided that $\mathop{\mathrm{Re}}\alpha>\max\{{(2-n)/p}-{1/(pp_{n})},{(2-n)/p}-(p-2)/[pp_{n}(p_{n}-2)]\}$ for $2\leq p\leq\infty$, with $p_{n}=2(n+1)/(n-1)$ for $n\geq 3$ and $p_{n}=4$ for $n=2$.
@@ 1.1. Background
provided that $\alpha>1-n+n/p$ when $1<p\leq 2;$ or $\alpha>(2-n)/p$ when $2\leq p\leq\infty.$ The above maximal theorem tells us that when $\alpha=0$ and $n\geq 3$, the operator $\mathfrak{M}$} is bounded on $L^{p}({\mathbb{R}^{n}})$ for p>n/(n1)p>n/(n-1). This range of $p$ is sharp. As is pointed out in [24, 25], no such result can hold for $p\leq n/(n-1)$ if $n\geq 2$.
@@
(1.5) $\|{\mathfrak{M}}^{\alpha}f\|_{L^{p}({\mathbb{R}^{2}})}\leq C\|f\|_{L^{p}({\mathbb{R}^{2}})},\ \ \ \mathop{\mathrm{Re}}\alpha>-\varepsilon(p),\ \ \ 2<p<\infty.$
@@
Note that in [17], Nowak, Roncal and Szarek [17] found sharp conditions for the operator ${\mathfrak{M}}^{\alpha}$ on radial functions to be bounded on $L^{p}_{\rm rad}({\mathbb{R}}^{n})$. They proved that for $n\geq 2$ and $\alpha>(1-n)/2$, ${\mathfrak{M}}^{\alpha}$ is bounded on $L^{p}_{\rm rad}({\mathbb{R}}^{n})$ if and only if $\alpha>1-n+n/p$ if $1<p\leq 2$; $\alpha>1/p+(1-n)/2$ if $2\leq p\leq 2n/(n-1)$; and $\alpha\geq(1-n)/p$ if p>2n/(n1)p>2n/(n-1).
@@ 1.2. Main results
The purpose of this article is to investigate the spherical maximal operators of (complex) order $\alpha$ on the $n$-dimensional hyperbolic space $\mathbb{H}^{n}$, which was introduced and studied by El Kohen ([13]). To do this, we denote the Minkowski metric on $\mathbb{R}^{n+1}$} by
$z,w]=z_{0}w_{0}-z_{1}w_{1}-\cdots-z_{n}w_{n},\ \ \ \ z=(z_{0},\dots,z_{n}),w=(w_{0},\dots,w_{n})\in{\mathbb{R}^{n+1}}.$
@@
and we equip this space with the Riemannian metric induced by the above Minkowski metric $\cdot$$\cdot$]. This Riemannian metric induces in turn a measure, by which we denote $\mathrm{d}z$. Define the spherical maximal operators by
@@
A natural question is for which $\alpha$ the spherical maximal opeators ${\mathfrak{m}}^{\alpha}$ are bounded on $L^{p}$. In [13, Theorem 3], El Kohen showed that for $n\geq 2$,
@@
(1.10) ${\rm Re}\,\alpha>1-n+{n\over p}\ \ \ \ {\rm when}\ 1<p\leq 2;$
or
(1.11) $\ \ {\rm Re}\,\alpha>{2-n\over p}\ \ \ \ \ \ \ \ {\rm when}\ 2<p\leq\infty.$
It was not until 2000 that Ionescu [10] showed that ${\mathfrak{m}}^{0}$ is bounded on $L^{p}(\mathbb{H}^{2})$ for $2<p\leq\infty$ by making use of the local smoothing estimates for Fourier integral operators ([20]). The above two admissible relations for $\alpha$ and $p$ are summarized when $n\geq 3$ in Figure 1.1 below, where the relations (1.10) and (1.11) correspond to the dashed segments $AB$ and $OB$, respectively.
@@
Figure $1.$ The $(1/p,\alpha)$-plot for $n\geq 3$. El Kohen’s result [13, Theorem 3] indicates that $\mathfrak{m}^{\alpha}$ is bounded on $L^{p}(\mathbb{H}^{n})$ if $(1/p,\alpha)$ is strictly above the dashed folded segments $OBA$. We extend the range of $(1/p,\alpha)$ in El Kohen’s result for $2<p<\infty$ to the range strictly above the dotted folded segments $OCB$. In addition, we showed that $\mathfrak{m}^{\alpha}$ is bounded on $L^{p}(\mathbb{H}^{n})$ only if $(1/p,\alpha)$ is on or above the solid folded segments ODBODB for $2<p<\infty$, or strictly above the dashed segment BABA for $1<p\leq 2$.
@@
(1.13) $\alpha\geq\max\left\{{1\over p}-{n-1\over 2},\ -{n-1\over p}\right\}\ \ \ \ {\rm when}\ 2<p\leq\infty.$
@@
Theorems  1.1 and  1.2 can be illustrated using Figure 1.1 above. The relations (1.12) and (1.13) in Theorem 1.1 correspond to the dashed segment $AB$ and the solid folded segments $ODB$, respectively. Theorem 1.2 indicates that compared to (1.11), there is an $1/(pp_{n})$-downwards extension of the range of $\mathop{\mathrm{Re}}\alpha$ so that (1.9) is valid for $p\geq p_{n}$, and there is a $(p-2)/[pp_{n}(p_{n}-2)]$-improvement for $2<p<p_{n}$. This is indicated in Figure 1.1 above by the dotted folded segments $OCB$.
@@ 2. Fourier analysis on the hyperbolic space
$z,w]=z_{0}w_{0}-z_{1}w_{1}-\cdots-z_{n}w_{n},\ \ \ \ \ z=(z_{0},\dots,z_{n}),\ w=(w_{0},\dots,w_{n})\in{\mathbb{R}^{n+1}}.$
@@
and equip this space with the Riemannian metric induced by the Minkowski metric $\cdot$$\cdot$]. This Riemannian metric induces in turn a measure, by which we denote $\mathrm{d}z$. We will further distinguish the point ${\bf 0}=(1,0,\dots,0)$.
@@
(notice that $z,b(\omega)]>0$ for $z\in\mathbb{H}^{n}$). They satisfy
@@
$\left|{\mathrm{d}^{k}\over\mathrm{d}\lambda^{k}}(\lambda^{-1}({\mathcal{H}}(\lambda))^{-1})\right|\leq C_{k}\,(1+|\lambda|)^{(n-1)/2-1-k}.$
@@
$\left|{\mathrm{d}^{k}\over\mathrm{d}\lambda^{k}}(\lambda\,{\mathcal{H}}^{\alpha}(\lambda))\right|\leq C_{\alpha,k}\,(1+|\lambda|)^{1-(n-1)/2-\mathop{\mathrm{Re}}\alpha-k}.$
@@
This is a consequence of Stirling’s formula ([27, p. 151, Example. (ii)]). ∎
@@
In [13, Theorem 1], El Kohen showed that for $\alpha\in\mathbb{C}$ with $\mathop{\mathrm{Re}}\alpha>(1-n)/2$ one can write
@@
$\times\left({\mathfrak{R}}^{\alpha}(\lambda,t)+t^{\alpha+(n-1)/2}\sum_{j=0}^{N}t^{2j}b^{\alpha}_{j}(t)\,\mathfrak{O}_{\alpha+(n-2)/2+j}(\lambda t)\right).$
@@
$\times F\!\left(-\alpha-{n-3\over 2},\alpha+{n-1\over 2};1-\mathrm{i}\lambda;{-1\over\mathrm{e}^{2t}-1}\right)$
@@
$\|\mathfrak{m}^{\alpha}(g_{j})\|_{p}$ $\geq$ $\left(\int_{\{z=(z_{0},z_{1},z^{\prime\prime})\in\mathbb{H}^{n}:\rho(z)\sim 1,|z^{\prime\prime}|\leq c_{2}2^{-j/2}\}}|M^{\alpha}_{\rho(z)}(g_{j})(z)|^{p}\,\mathrm{d}z\right)^{1/p}$
@@
Now we consider $z\in B_{\varepsilon/100}(\mathbf{0})$. Then it’s obvious that $\mathop{\mathrm{supp}}h_{\varepsilon}$ is contained in the interior of $B_{1}(z)$, which indicates that if $w\in\mathop{\mathrm{supp}}h_{\varepsilon}$, then we have 1d(z,w)>01-d(z,w)>0. Thus in view of (1.7) for such $z$ we can write
@@
where $a^{\alpha}_{2}(\lambda,t)\in S^{0}_{\lambda}$ and ${\mathcal{H}}^{\alpha}$ is as in (2.4). Set $\phi_{s}\in C^{\infty}_{c}(\mathbb{R})$, $0\leq\phi_{s}\leq 1$, $\phi_{s}\equiv 1$ on $[0,s+1]$ and $\phi_{s}\equiv 0$ outside (1,s+2)(-1,s+2). The convolution kernel $K^{\alpha}_{t}(r)$ of $M^{\alpha}_{t}$ can be written as the following:
@@
$\times\int_{\mathbb{R}}\beta_{j}(\lambda)\,{\mathcal{H}}^{\alpha}(\lambda)\left({\mathrm{e}^{\mathrm{i}\lambda(t+r)}\over{\mathcal{H}}(-\lambda)}a_{2}^{0}(\lambda,r)+{\mathrm{e}^{\mathrm{i}\lambda(t-r)}\over{\mathcal{H}}(\lambda)}a_{2}^{0}(-\lambda,r)\right)a_{2}^{\alpha}(\lambda,t)\,\mathrm{d}\lambda$
@@
$\times\int_{0}^{\infty}\beta_{j}(\lambda)\left(\mathrm{e}^{\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(\lambda)\,a^{\alpha}_{2}(\lambda,t)+\mathrm{e}^{-\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(-\lambda)\,a^{\alpha}_{2}(-\lambda,t)\right)\varphi_{\lambda}(r)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda.$
@@
Now we assume $j\geq 1$. We take an even cutoff function $\gamma\in C^{\infty}_{c}(\mathbb{R})$ with $\gamma\equiv 1$ on [1,1][-1,1] and $\gamma\equiv 0$ outside (2,2)(-2,2). By Lemma 2.4(ii), we write
@@
$\times\int_{0}^{\infty}\beta_{j}(\lambda)\,\gamma(\lambda r)\left(\mathrm{e}^{\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(\lambda)\,a^{\alpha}_{2}(\lambda,t)+\mathrm{e}^{-\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(-\lambda)\,a^{\alpha}_{2}(-\lambda,t)\right)\varphi_{\lambda}(r)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda,$
@@
$\times\int_{\mathbb{R}}\beta_{j}(\lambda)\,(1-\gamma(\lambda r))\left(\mathrm{e}^{\mathrm{i}\lambda(t+r)}a^{0,N}_{1}(\lambda,r)+\mathrm{e}^{\mathrm{i}\lambda(t-r)}a^{0,N}_{1}(-\lambda,r)\right){\mathcal{H}}^{\alpha}(\lambda)\,a^{\alpha}_{2}(\lambda,t)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda,$
@@
$\times\int_{0}^{\infty}\beta_{j}(\lambda)\,(1-\gamma(\lambda r))\left(\mathrm{e}^{\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(\lambda)\,a^{\alpha}_{2}(\lambda,t)+\mathrm{e}^{-\mathrm{i}\lambda t}{\mathcal{H}}^{\alpha}(-\lambda)\,a^{\alpha}_{2}(-\lambda,t)\right)E^{\alpha,N}(\lambda,r)\,|{\mathcal{H}}(\lambda)|^{-2}\,\mathrm{d}\lambda.$
@@
$\sup_{z\in\mathbb{H}^{n}}\left\|\sup_{t\in[T,T+1]}|\mathscr{E}^{\alpha}_{t,j}(d(z,\cdot))|\right\|_{1}+\sup_{w\in\mathbb{H}^{n}}\left\|\sup_{t\in[T,T+1]}|\mathscr{E}^{\alpha}_{t,j}(d(\cdot,w))|\right\|_{1}\leq C_{\alpha,\sigma}2^{-(\mathop{\mathrm{Re}}\alpha+\sigma)j}\mathrm{e}^{-(n-1)T/2}$
@@
Let us now prove (4.22). To do this, we define another cutoff function $\widetilde{\phi}_{s}\in C_{c}^{\infty}(\mathbb{R})$ by $\widetilde{\phi}_{s}\equiv\phi_{s}$ on $s,\infty)$ and $\widetilde{\phi}_{s}\equiv 0$ on $(-\infty,s-1]$. Note that in view of (4.12), we can write
@@
In view of (4.25), we set the function $\Phi((v,u),(v^{\prime},u^{\prime}))$ in (4.4) to be $d((v,u),(v^{\prime},u^{\prime}-T_{0}))$ and the symbol $a(\lambda,t,(v,u),(v^{\prime},u^{\prime}))$ in (4.4) to be $b^{\alpha}(\lambda,t,d((v,u),(v^{\prime},u^{\prime}-T_{0})))$. Notice that the problem in $(v,u),(v^{\prime},u^{\prime})]-$coordinates is no longer degenerate. Indeed, on the one hand, by (4.29) the Monge-Ampere matrix of $\Phi$ turns into
@@
which is (4.2). Here $\nabla$ is the gradient and $|\cdot|_{g}$ is the norm both induced by the Riemannian metric on $\mathbb{H}^{n}$. Also, in view of (4.26) and (4.29), $b^{\alpha}(\lambda,t,d((v,u),(v^{\prime},u^{\prime}-T_{0})))$ satisfies (4.3) with $m$ replaced by $\mathop{\mathrm{Re}}\alpha$. Hence by (4.7) in Proposition 4.3, we can write
@@
Now we consider (4.33) for the case $0<t\leq 1$. Let $\gamma$ be the cutoff function in the proof of Lemma 4.6. We define another cutoff function $\widetilde{\gamma}\in C_{c}^{\infty}(\mathbb{R}_{+})$ with $\widetilde{\gamma}\equiv 1$ on $1/2,\infty)$ and $\widetilde{\gamma}\equiv 0$ on $[0,1/4]$. Let $K^{\alpha}_{t}(r)$ be the convolution kernel of $M^{\alpha}_{t}$. Note that in view of (1.7), $\mathop{\mathrm{supp}}K^{\alpha}_{t}(r)\subset\{r:0\leq r\leq t\}$. So we can write
@@
Acknowledgements: The authors were supported by National Key R&\&D Program of China 2022YFA1005700. P. Chen was supported by NNSF of China 12171489. M. Shen was supported by China Postdoctoral Science Foundation 2024M761509. L. Yan was supported by NNSF of China 12571111. The authors thank X. Chen, N. Liu, L. Roncal, L. Song and H.-W. Zhang for helpful discussions and suggestions.